# Diophantine Graphs

### v1

@@ 4.3 More Necessary and Sufficient Conditions for Equality of Degrees
The following theorem is a generalization of one direction of Theorems 4.3.
@@
Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f$ and $u,v\in V(D_{n})$ such that $\deg(u)<n-1$, $\deg(v)<n-1$. If $\omega\big{(}f^{*}(u)\big{)}=\omega\big{(}f^{*}(v)\big{)}$ and for every prime number $p\mid f^{*}(u)$, there exits a prime number $q\mid f^{*}(v)$ such that $\left\lfloor\frac{n}{p^{\acute{v}_{p}(n)}}\right\rfloor=\left\lfloor\frac{n}{q^{\acute{v}_{q}(n)}}\right\rfloor$, then $\deg(u)=\deg(v)$, where $w(n)$ is the number of prime factors of $n$.
@@
Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f$ and $u,v\in V(D_{n})$ such that
$\deg(u)<n-1,\quad\deg(v)<n-1\quad\mbox{and}\quad\omega\big{(}f^{*}(u)\big{)}=\omega\big{(}f^{*}(v)\big{)}.$
Therefore, $\omega\big{(}f^{*}(u)\big{)}=\omega\big{(}f^{*}(v)\big{)}>0.$ Hence, there is a prime number $p$ such that $p\mid f^{*}(u)$. Consequently, our hypothesis implies that there is a prime number $q\mid f^{*}(v)$ such that
@@
Therefore, using Lemmas 2.2 part $(i)$, $p^{\acute{v}_{p}(n)}\mid f(u)$ and $q^{\acute{v}_{q}(n)}\mid f(v).$ Then, using the fundamental theorem of arithmetic, there exist $\grave{u},\grave{v}\in V(D_{n})$ such that
$f(\grave{u}):=\prod\limits_{p\mid f^{*}(u)}p^{\acute{v}_{p}(n)}\ \Big{|}\ f(u)\quad\mbox{and}\quad f(\grave{v}):=\prod\limits_{q\mid f^{*}(v)}q^{\acute{v}_{q}(n)}\ \Big{|}\ f(v).$
By definition of $f(\grave{u})$ and $f(\grave{v})$, we have $f^{*}(u),f^{*}(\grave{u})$ have the same prime factors and $f^{*}(v),f^{*}(\grave{v})$ have the same prime factors. Since $\omega\big{(}f^{*}(u)\big{)}=\omega\big{(}f^{*}(v)\big{)},$ therefore
$\omega\big{(}f^{*}(\grave{u})\big{)}=\omega\big{(}f^{*}(u)\big{)}=\omega\big{(}f^{*}(v)\big{)}=\omega\big{(}f^{*}(\grave{v})\big{)}.$ (9)
Moreover, for every prime number $p$, we have
$p^{\acute{v}_{p}(n)}\mid f(u)\ \mbox{if and only if}\ p^{\acute{v}_{p}(n)}\mid f(\grave{u}).$
@@
$\deg(v)=\deg(\grave{v}).$ (11)
Let $d=(f(\grave{u}),f(\grave{v}))$. Thus, Lemmas 2.2 part $(ii)$ implies that
$\frac{d}{(d,n)}\ \bigg{|}\ f^{*}(\grave{u}),f^{*}(\grave{v}),$ (12)
and hence
$\omega\left(\frac{d}{(d,n)}\right)\leq\omega\big{(}f^{*}(\grave{u})\big{)}=\omega\big{(}f^{*}(\grave{v})\big{)}.$
Let us consider the following cases.
Case 1: Let $\omega\big{(}f^{*}(\grave{u})\big{)}=\omega\big{(}f^{*}(\grave{v})\big{)}=\omega\left(\frac{d}{(d,n)}\right).$ Then, using equation (12), $\frac{d}{(d,n)}$, $f^{*}(\grave{u})$ and $f^{*}(\grave{v})$ have the same prime factors. A similar argument used to prove equation (10) implies that
$\deg(\grave{u})=\deg(\grave{v}).$
Hence, using equations (10) and (11), we get
$\deg(u)=\deg(v).$
Case 2: Let $\omega\big{(}f^{*}(\grave{u})\big{)}=\omega\big{(}f^{*}(\grave{v})\big{)}=\omega\left(\frac{d}{(d,n)}\right)+1=r.$ Then we get $f^{*}(\grave{u})$ and $f^{*}(\grave{v})$ differ in exactly one prime factor. Therefore, using the fundamental theorem of arithmetic, we obtain the following two equations
@@
$\begin{split}\deg(\grave{u})=n-\sum_{1\leq i\leq r-1}\left\lfloor\frac{n}{h_{i}^{\acute{v}_{h_{i}}(n)}}\right\rfloor&-\left\lfloor\frac{n}{p^{\acute{v}_{p}(n)}}\right\rfloor+\sum_{1\leq i<j\leq r-1}\left\lfloor\frac{n}{h_{i}^{\acute{v}_{h_{i}}(n)}h_{j}^{\acute{v}_{h_{j}}(n)}}\right\rfloor+\sum_{1\leq i\leq r-1}\left\lfloor\frac{n}{h_{i}^{\acute{v}_{h_{i}}(n)}p^{\acute{v}_{p}(n)}}\right\rfloor-\\ &-\dots+(-1)^{r}\left\lfloor\frac{n}{\prod\limits_{1\leq i\leq r-1}h_{i}^{\acute{v}_{h_{i}}(n)}p^{\acute{v}_{p}(n)}}\right\rfloor,\end{split}$ (13)
@@
$\begin{split}\deg(\grave{v})=n-\sum_{1\leq i\leq r-1}\left\lfloor\frac{n}{h_{i}^{\acute{v}_{h_{i}}(n)}}\right\rfloor&-\left\lfloor\frac{n}{q^{\acute{v}_{q}(n)}}\right\rfloor+\sum_{1\leq i<j\leq r-1}\left\lfloor\frac{n}{h_{i}^{\acute{v}_{h_{i}}(n)}h_{j}^{\acute{v}_{h_{j}}(n)}}\right\rfloor+\sum_{1\leq i\leq r-1}\left\lfloor\frac{n}{h_{i}^{\acute{v}_{h_{i}}(n)}q^{\acute{v}_{q}(n)}}\right\rfloor-\\ &-\dots+(-1)^{r}\left\lfloor\frac{n}{\prod\limits_{1\leq i\leq r-1}h_{i}^{\acute{v}_{h_{i}}(n)}q^{\acute{v}_{q}(n)}}\right\rfloor.\end{split}$ (14)
Using Lemma 2.6 and equations (8), (13) and (14), we get
$\deg(\grave{u})=\deg(\grave{v}),$
and hence
$\deg(u)=\deg(v).$
Case 3: Let $\omega\big{(}f^{*}(\grave{u})\big{)}=\omega\big{(}f^{*}(\grave{v})\big{)}=\omega\left(\frac{d}{(d,n)}\right)+2=r.$ Then $f^{*}(\grave{u})$ and $f^{*}(\grave{v})$ differ in exactly two prime factors. Therefore, using the fundamental theorem of arithmetic, we have the following two equations
@@
Define
$a:=p_{1}^{\acute{v}_{p_{1}}(n)}p_{2}^{\acute{v}_{p_{2}}(n)}\quad\mbox{and}\quad b:=q_{1}^{\acute{v}_{q_{1}}(n)}q_{2}^{\acute{v}_{q_{1}}(n)}.$
Since
$\frac{a}{(a,n)}=p_{1}p_{2}\quad\mbox{and}\quad\frac{b}{(b,n)}=q_{1}q_{2},$
therefore Lemma 2.5 implies that $a>n\ \mbox{or}\ b>n.$ Hence,
$f(\grave{u})\geq a>n\quad\mbox{or}\quad f(\grave{v})\geq b>n,$
which is absurd. The cases $\omega\big{(}f^{*}(\grave{u})\big{)}=\omega\big{(}f^{*}(\grave{v})\big{)}=\omega\left(\frac{d}{(d,n)}\right)+i$, for every $i\geq 3$ can be treated in a very similar way of case $3$ which are absurd. Hence the result follows. ∎
@@
In Example 4.2, let $f(u)=14$ and $f(v)=16$, where $u,v\in V(D_{20})$, so $f^{*}(u)=7$ and $f^{*}(v)=8$. Then, applying Theorem 4.5, we have $\omega\left(7\right)=\omega\left(8\right)=1$ and $\left\lfloor\frac{20}{7}\right\rfloor=\left\lfloor\frac{20}{8}\right\rfloor=2.$ Hence, $\deg(u)=\deg(v)=18$, (by applying Lemma 4.1).
The converse of Theorem 4.5 does not hold. For instance, let $u,v\in V(D_{23})$ such that $f(u)=10$, $f(v)=14$. By Theorem 4.2, we have $deg(u)=deg(v)=10$. However, it can be verified that $\left\lfloor\frac{23}{5}\right\rfloor\neq\left\lfloor\frac{23}{7}\right\rfloor$, which demonstrates that the equality of degrees does not imply the equality of values of the floor function.
@@
Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f$. Let $u,v\in V(D_{n})$ such that $f^{*}(u)>1$, $f^{*}(v)>1$. If $f^{*}(u),f^{*}(v)$ have the same prime factors, then $\deg(u)=\deg(v)$.
@@
Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f$. Let $u,v\in V(D_{n})$ such that $f(v)=tf(u)$ for some $t\geq 1$. If $t\mid n$ and $(t,f(u))=1$, then $\deg(u)=\deg(v)$.
@@
Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f$. Let $u,v\in V(D_{n})$ such that $f(v)=tf(u)$ for some $t\geq 1$, $t\mid n$ and $(t,f(u))=1$. The reduced labels $f^{*}(v),f^{*}(u)$ satisfy the following equalities.
@@
Therefore, the two reduced labels $f^{*}(u),f^{*}(v)$ have the same primes factors or each of them is equal to 1. Hence, using Corollary 4.3 and Theorem 4.1, the result follows. ∎
The following two examples shows the converse of Corollary 4.3 is not true,
@@
Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f$. Let $u,v\in V(D_{n})$ such that $f(u)\neq f(v)$, $f(u)\mid f(v)$, and $f(v)=p^{\acute{v}_{p}(n)}$, $\frac{n}{2}<f(v)<n$, where $p$ is a prime number, using Theorem 4.1, we have
$\deg(u)=n-1=\deg(v).$
However, $f^{*}(u)=\frac{f(u)}{(f(u),n)}=1$ and $f^{*}(v)=\frac{p^{\acute{v}_{p}(n)}}{\left(p^{\acute{v}_{p}(n)},n\right)}=p$ have not the same prime factors.
@@
Suppose $D_{n}$ is the maximal Diophantine graph of even order $n$ with a labeling $f$. Let $u,v\in V(D_{n})$ such that $f(u)=2p^{\acute{v}_{p}(n)}$ and $f(v)=2q^{\acute{v}_{q}(n)}$, $\frac{n}{3}<p^{\acute{v}_{p}(n)},q^{\acute{v}_{q}(n)}<\frac{n}{2}$, where $p,q$ are distinguish prime numbers, this means $f(u)\nmid f(v)$, using Lemma 4.1 and Lemmas 2.2 part $(i)$, we have
@@
Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f$. Let $u,v\in V(D_{n})$ such that $f(u)\mid f(v)$, $f(v)$ is not a prime power number and $f^{*}(u)>1$. If $\deg(u)=\deg(v)$, then $f^{*}(u),f^{*}(v)$ have the same prime factors.
@@
Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f$. Let $u,v\in V(D_{n})$ such that $f(u)\mid f(v)$, $f(v)$ is not a prime power number, $\deg(u)=\deg(v)$ and $f^{*}(u)>1$. Then, using Lemmas 2.2 part $(ii)$, we have $f^{*}(u)\mid f^{*}(v)$. Consequently, $f^{*}(v)>1$ and also any prime factor of $f^{*}(u)$ is a prime factor of $f^{*}(v)$. Let there exists a prime number $p$ such that $p\mid f^{*}(v)$. Then, using Lemmas 2.2 part $(i)$, $p^{\acute{v}_{p}(n)}\mid f(v)$, i.e., $f(v)\in M_{p^{\acute{v}_{p}(n)}}$. Let a vertex $s\in V(D_{n})$ such that $f(s)=p^{\acute{v}_{p}(n)}\in M_{p^{\acute{v}_{p}(n)}}.$ Since $f(v)$ is not a prime power number, we get $s\neq v$. Using Lemma 3.1, we have $(f(v),f(s))\nmid n$. Therefore, $s\notin N(v)$. Consequently, using Lemma 4.3, we obtain $s\notin N(u)$. Then we get $\left(f(u),p^{\acute{v}_{p}(n)}\right)\nmid n$. Thus, $p^{\acute{v}_{p}(n)}\mid f(u)$. Consequently, using Lemmas 2.2 part $(i)$, $p\mid f^{*}(u)$. Therefore, for every prime number $p$,
@@
Hence, the two reduced labels $f^{*}(u),f^{*}(v)$ have the same prime factors. ∎
Future Work
Give two efficient algorithms that one determines the maximal Diophantine graphs and anther determines whether a given graph is Diophantine or not. Furthermore, exploring different variants of Diophantine labeling can further enrich the theory and uncover new connections between graph theory and number theory.
Declaration
The authors contributed equally to the manuscripts and approved the submitted scripts. We declare that we have no conflicts of interests. We confirm that this work is original and has not been published elsewhere, nor is it currently under consideration for publication elsewhere.

### v2

@@ 4.3 More Necessary and Sufficient Conditions for Equality of Degrees
The following theorem is a generalization of one direction of Theorem 4.4.
@@
Let $u,v$ be two vertices in $D_{n}$ with $\deg(u)<n-1$ and $\deg(v)<n-1$. If $\omega\big(f^{*}(u)\big)=\omega\big(f^{*}(v)\big)$ and for every prime number $p\mid f^{*}(u)$, there exists a prime number $q\mid f^{*}(v)$ such that $\left\lfloor\frac{n}{p^{\acute{v}_{p}(n)}}\right\rfloor=\left\lfloor\frac{n}{q^{\acute{v}_{q}(n)}}\right\rfloor$, then $\deg(u)=\deg(v)$, where $w(n)$ is the number of distinct prime factors of $n$.
@@
Let $u,v$ be two vertices in $D_{n}$ such that $\deg(u)<n-1$, $\deg(v)<n-1$ and
$\omega\big(f^{*}(u)\big)=\omega\big(f^{*}(v)\big).$ (8)
Then, using Lemma 4.4, there exists a prime number $p$ such that $p^{\acute{v}_{p}(n)}\mid f(u).$ Therefore, using Lemma 2.1, part i., there exists a prime number $p$ such that $p\mid f^{*}(u)$. Consequently, It follows from our assumption that there exists a prime number $q$ such that $q\mid f^{*}(v)$ and
@@
Moreover, $\omega\big(f^{*}(u)\big)=\omega\big(f^{*}(v)\big)>0.$ Since $p^{\acute{v}_{p}(n)}\mid f(u)$ and $q^{\acute{v}_{q}(n)}\mid f(v)$ and using Theorem 1.1, therefore there exist $\grave{u},\grave{v}\in V(D_{n})$ such that
$f(\grave{u}):=\prod\limits_{p\mid f^{*}(u)}p^{\acute{v}_{p}(n)}\ \bigg|\ f(u)\quad\mbox{and}\quad f(\grave{v}):=\prod\limits_{q\mid f^{*}(v)}q^{\acute{v}_{q}(n)}\ \bigg|\ f(v).$
[…]
$).\omega\big(f^{*}(\grave{u})\big)=\omega\big(f^{*}(u)\big)=\omega\big(f^{*}(v)\big)=\omega\big(f^{*}(\grave{v}$$\big)$. (10)
Furthermore, for every prime number $p$, we have
[…]
@@
[…]
Let $)d=(f(\grave{u}),f(\grave{v}$)). Thus, Lemmas 2.1, part ii.ii., implies that
[…]
@@
$\begin{split}\deg(\grave{u})=n&-\sum_{1\leq i\leq r-1}\left\lfloor\frac{n}{h_{i}^{\acute{v}_{h_{i}}(n)}}\right\rfloor-\left\lfloor\frac{n}{p^{\acute{v}_{p}(n)}}\right\rfloor+\sum_{1\leq i<j\leq r-1}\left\lfloor\frac{n}{h_{i}^{\acute{v}_{h_{i}}(n)}h_{j}^{\acute{v}_{h_{j}}(n)}}\right\rfloor\\ &+\sum_{1\leq i\leq r-1}\left\lfloor\frac{n}{h_{i}^{\acute{v}_{h_{i}}(n)}p^{\acute{v}_{p}(n)}}\right\rfloor-\dots+(-1)^{r}\left\lfloor\frac{n}{\prod\limits_{1\leq i\leq r-1}h_{i}^{\acute{v}_{h_{i}}(n)}p^{\acute{v}_{p}(n)}}\right\rfloor,\end{split}$ (14)
@@
$\begin{split}\deg(\grave{v})=n&-\sum_{1\leq i\leq r-1}\left\lfloor\frac{n}{h_{i}^{\acute{v}_{h_{i}}(n)}}\right\rfloor-\left\lfloor\frac{n}{q^{\acute{v}_{q}(n)}}\right\rfloor+\sum_{1\leq i<j\leq r-1}\left\lfloor\frac{n}{h_{i}^{\acute{v}_{h_{i}}(n)}h_{j}^{\acute{v}_{h_{j}}(n)}}\right\rfloor\\ &+\sum_{1\leq i\leq r-1}\left\lfloor\frac{n}{h_{i}^{\acute{v}_{h_{i}}(n)}q^{\acute{v}_{q}(n)}}\right\rfloor-\dots+(-1)^{r}\left\lfloor\frac{n}{\prod\limits_{1\leq i\leq r-1}h_{i}^{\acute{v}_{h_{i}}(n)}q^{\acute{v}_{q}(n)}}\right\rfloor.\end{split}$ (15)
[…]
@@
Define $a:=p_{1}^{\acute{v}_{p_{1}}(n)}p_{2}^{\acute{v}_{p_{2}}(n)}\ \mbox{and}\ b:=q_{1}^{\acute{v}_{q_{1}}(n)}q_{2}^{\acute{v}_{q_{1}}(n)}.$ Since $\frac{a}{(a,n)}=p_{1}p_{2}\ \mbox{and}\ \frac{b}{(b,n)}=q_{1}q_{2},$ therefore Lemma 2.4 implies that $a>n\ \mbox{or}\ b>n.$ Hence, $f(\grave{u})\geq a>n\ \mbox{or}\ f(\grave{v})\geq b>n,$ which is absurd.
The cases $\omega\big(f^{*}(\grave{u})\big)=\omega\big(f^{*}(\grave{v})\big)=\omega\left(\frac{d}{(d,n)}\right)+i$, for every $i>3$ can be treated in a very similar way of case $3$ which are absurd. Hence the proof follows. ∎
@@
In $D_{20}$, let $f(u)=14$ and $f(v)=16$, where $u,v\in V(D_{20})$, so $f^{*}(u)=7$ and $f^{*}(v)=8$. Then, applying Theorem 4.6, we have $\omega\left(7\right)=\omega\left(8\right)=1$ and $\left\lfloor\frac{20}{7}\right\rfloor=\left\lfloor\frac{20}{8}\right\rfloor=2.$ Hence, $\deg(u)=\deg(v)=18$, (by applying Lemma 4.1).
The converse of Theorem 4.6 is not true. For instance, in $D_{23}$, let $f(u)=10$, $f(v)=14$, where $u,v\in V(D_{23})$. Then, by Theorem 4.3, we have $deg(u)=13=deg(v)$. However, it can be verified that $\left\lfloor\frac{23}{5}\right\rfloor\neq\left\lfloor\frac{23}{7}\right\rfloor$.
@@
Let $u,v$ be two vertices in $D_{n}$ with $f^{*}(u)>1$, $f^{*}(v)>1$. If $f^{*}(u),f^{*}(v)$ have the same prime factors, then $\deg(u)=\deg(v)$.
@@
Let $u,v$ be two vertices in $D_{n}$ such that $f(v)=tf(u)$ for some $t\in\mathbb{Z^{+}}$. If $(t,f(u))=1$ and $t\mid n$, then $\deg(u)=\deg(v)$.
@@
Let $u,v$ be two vertices in $D_{n}$ such that $f(v)=tf(u)$ for some $t\in\mathbb{Z^{+}}$, $(t,f(u))=1$ and $t\mid n$. The reduced labels $f^{*}(v),f^{*}(u)$ satisfy the following equalities.
@@
Therefore, $f^{*}(u),f^{*}(v)$ have the same prime factors or $f^{*}(u)=f^{*}(v)=1$. Hence, using Corollary 4.3 and Theorem 4.1, the proof follows. ∎
The following two examples show that the converse of Corollary 4.3 is not true.
@@
Let $u,v$ be two different vertices in $D_{n}$ such that $f(u)\mid f(v)$ and $f(v)=p^{\acute{v}_{p}(n)}$, $\frac{n}{2}<f(v)<n$, where $p$ is a prime number. Then, using Theorem 4.1, we have $\deg(u)=n-1=\deg(v).$ However, $f^{*}(u)=\frac{f(u)}{(f(u),n)}=1$ and $f^{*}(v)=\frac{p^{\acute{v}_{p}(n)}}{\left(p^{\acute{v}_{p}(n)},n\right)}=p$ have not the same prime factors.
@@
Let $u,v$ be two vertices in $D_{n}$ such that $f(u)=2p^{\acute{v}_{p}(n)}$ and $f(v)=2q^{\acute{v}_{q}(n)}$, $\frac{n}{3}<p^{\acute{v}_{p}(n)},q^{\acute{v}_{q}(n)}<\frac{n}{2}$, where $p,q$ are distinguish prime numbers, this means $f(u)\nmid f(v)$. Then, using Lemma 4.1 and Lemmas 2.1, part $i.$, we have
@@
Let $u,v$ be two vertices in $D_{n}$ such that $f(u)\mid f(v)$, $f(v)$ is not a prime power and $f^{*}(u)>1$. If $\deg(u)=\deg(v)$, then $f^{*}(u),f^{*}(v)$ have the same prime factors.
@@
Let $u,v$ be two vertices in $D_{n}$ such that $f(u)\mid f(v)$, $f(v)$ is not a prime power, $\deg(u)=\deg(v)$ and $f^{*}(u)>1$. Then, using Lemmas 2.1, part ii.ii., we have $f^{*}(u)\mid f^{*}(v)$. Consequently, $f^{*}(v)>1$ and also any prime factor of $f^{*}(u)$ is a prime factor of $f^{*}(v)$. Let there exists a prime number $p$ such that $p\mid f^{*}(v)$. Then, using Lemmas 2.1, part $i.$, $p^{\acute{v}_{p}(n)}\mid f(v)$, i.e., $f(v)\in M_{p^{\acute{v}_{p}(n)}}$. Let a vertex $s\in V(D_{n})$ such that $f(s)=p^{\acute{v}_{p}(n)}\in M_{p^{\acute{v}_{p}(n)}}.$ Since $f(v)$ is divisible by at least two distinct prime numbers, we get $s\neq v$. Using Lemma 3.1, we have $(f(v),f(s))\nmid n$. Therefore, $s\notin N(v)-\{u\}$. Consequently, using Lemma 4.3, we obtain $s\notin N(u)-\{v\}$. Then we get $\left(f(u),f(s)\right)\nmid n$. Thus, $f(s)=p^{\acute{v}_{p}(n)}\mid f(u)$. Consequently, using Lemmas 2.1, part $i.$, $p\mid f^{*}(u)$. Therefore, for every prime number $p$,
@@
Hence, the proof follows. ∎
5 Conclusions
This manuscript introduces the concept of Diophantine labelling and establishes a connection between Diophantine graphs and prime graphs. We study further define maximal Diophantine graphs and derives explicit formulas for their edge count and vertex degrees. Additionally, we provide necessary and sufficient conditions for the equality of vertex degrees. We also utilize number-theoretic concepts, such as p-adic valuation, to analyze vertex labels and the properties of Diophantine graphs. Although Diophantine labelling is still primarily theoretical, we think that Diophantine labelling sits at an intersection of graph theory, number theory and combinatorics. Furthermore, exploring different variants of Diophantine labelling can further enrich the theory and uncover new connections between these mathematical domains. As research advances, we expect that Diophantine labelling has more concrete applications in the future. Moreover, the algorithmic approach is promising in many directions, including searching for an efficient algorithm for Diophantine graph and finding maximal Diophantine graphs.
Acknowledgment
The author would like to extend their gratitude to the referees for their valuable feedback, insightful comments, and constructive suggestions, which significantly contributed to improving the quality and clarity of this work.